OpenStudy (anonymous):

Find the equation of the line that contains the point (14, -9) and has slope -3 - 7

OpenStudy (anonymous):

lol, i don't know how people get stuck in such simple problems. i advice, go read the concept and formulae once again. its no mind game, its just memory game. lol

OpenStudy (anonymous):

sorry, i was replaying to s

OpenStudy (anonymous):

sorry, i was replaying to some other thread and by mistake posted it here. see, the equation of line is y = m.x + c where m = slope and c = intercept now you know the slope i.e. -3 the only thing you gotta know is the intercept now as the line passes through point (14, -9) so it satisfies the equation. therefore, -9 = -3 (14) + c => c = -9 + 3 (14) => c = 51 therefore the required equation is y = -3x +51